Given the conic section -3.6x2+1.6y2+72x+6.4y = 58/5

a) State the nature of the conic section

b) Find the characteristic elements of the conic section

c) Sketch the conic section indicating all the elements.


1
Expert's answer
2021-09-28T15:01:50-0400

a)


"-3.6x^2+1.6y^2+72x+6.4y = 58\/5"

"-3.6(x^2-20x+100)+360+1.6(y^2+4y+4)-6.4=11.6"

"3.6(x-10)^2-1.6(y+2)^2=342"

"\dfrac{(x-10)^2}{95}-\dfrac{(y+2)^2}{213.75}=1"

A conic is a hyperbola. Standard form


"\dfrac{(x-10)^2}{95}-\dfrac{(y+2)^2}{213.75}=1"

Horizontal hyperbola.

b)

"h=10, k=-2, a=\sqrt{95}, b=\sqrt{213.75}=1.5\sqrt{95}"

"c^2=a^2+b^2=95+213.75=308.75, c=\sqrt{308.75}"

Center: "(h, k)=(10,-2)"

Vertices: "(h\pm a,k), (10-\sqrt{95}, -2), (10+\sqrt{95}, -2)"

Covertices: "(h, k\pm b), (10, -2-1.5\sqrt{95}), (10, -2+1.5\sqrt{95})"

Foci: "(h\pm c,k), (10-\sqrt{308.75}, -2), (10+\sqrt{308.75}, -2)"

The equations of the asymptotes are 

"y=\dfrac{b}{a}(x-h)+k,y=-\dfrac{b}{a}(x-h)+k""y=1.5(x-10)-2,y=-1.5(x-10)-2"

"x=0,\dfrac{(0-10)^2}{95}-\dfrac{(y+2)^2}{213.75}=1"


"(y+2)^2=11.24"

"y=-2\pm1.5\sqrt{5}"

"(0, -2-1.5\sqrt{5}), (0, -2+1.5\sqrt{5})"



"y=0,\dfrac{(x-10)^2}{95}-\dfrac{(0+2)^2}{213.75}=1"


"(x-10)^2=\dfrac{839}{9}"

"x=10\pm\dfrac{\sqrt{839}}{3}"

"(10-\dfrac{\sqrt{871}}{3},0), (10+\dfrac{\sqrt{871}}{3}, 0)"

c)